OpenStudy (anonymous):

MEDAL/FAN4FAN HELP Graph the function. y=lx+2l-3 I know there's an easy way to figure out the placement and slope just from using the function above.. but I'm not exactly sure! I will post the options.

OpenStudy (anonymous):

I got this.

OpenStudy (jdoe0001):

well, the one thing you'd need to know is that |x+2| stand for +1(x+2) and -1(x+2) so y=lx+2l-3 is really meant to be y=+1(x+2)-3 AND y=-1(x+2)-3 so, graph those two

OpenStudy (anonymous):

I posted the options above, would it be D then?

OpenStudy (jdoe0001):

dunno, graph them

OpenStudy (anonymous):

I tried.

OpenStudy (jdoe0001):

well, so, match your graph against the given choices

OpenStudy (anonymous):

I just want someone to check my answer then because once I submit I can't change it. I got D.

OpenStudy (jdoe0001):

which means he didn't graph it :/

OpenStudy (anonymous):

I am a she, and I did graph it thank you. But it is distorted so I must have graphed it incorrectly.

OpenStudy (anonymous):

You don't need to talk down to me, I am just trying to find some help. Explain graphing then perhaps?

OpenStudy (jdoe0001):

heheh, ok, She did graph it then, it's just 2 lines you know, so pick 2 values for "x" and plot those over a straight line, do the same for the 2nd equation

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

y=+1(x+2)-3 y=-1(x+2)-3 These right ^ I still don't quite understand I have to be doing it wrong... I'm more for equations then graphing hahaha

OpenStudy (jdoe0001):

<wait with anxiety> <biting nails>

OpenStudy (anonymous):

input for x, like 1 and 2 or whatever right

OpenStudy (anonymous):

then what

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

what would y equal and x equal to be able to graph it?

OpenStudy (anonymous):

first one y=1 if I use 1 as x

OpenStudy (anonymous):

second y=-1

OpenStudy (jdoe0001):

well, the 1st equation is y = +(x+2)-3 so make x = 1 y = +(1+2)-3 => 0

OpenStudy (anonymous):

so it sits on the x axis?

OpenStudy (anonymous):

hahahah I still don't understanddd

OpenStudy (jdoe0001):

well, it gives you ... yes, it "touches" the y-axis

OpenStudy (jdoe0001):

wait no, the x-axis

OpenStudy (jdoe0001):

so, it touches the x-axis, at x = 1

OpenStudy (anonymous):

it's A...?

OpenStudy (anonymous):

I think I got the gist.

OpenStudy (anonymous):

is it A or not hahah

OpenStudy (jdoe0001):

hehe

OpenStudy (anonymous):

because in A it correlates with the graphing right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

sweet. I need more help can you do like speed step by step if I get the answers wrong? just a few more questions I'm checkingg

OpenStudy (jdoe0001):

well, is easier if you post in channel, that way we can all see it and help and revise earch other :)

OpenStudy (anonymous):

channel?

OpenStudy (jdoe0001):

so-called main channel, so close this one, and open/ask a new one :)